OpenStudy (anonymous):

Let f(x) = 3x. Find f( 2)

OpenStudy (anonymous):

put 2 instead of x

OpenStudy (diyadiya):

f(x)=3x f(2)3*2=6

OpenStudy (anonymous):

f(2)=3*2=6